Operations Report – March 1st

Crew 206 Operations Report 01 – MAR – 2019

SOL: 5

Name of person filing report: Gaspard Thieulin

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and
alternator need replacement on generator. A contractor has
been arranged to make these repairs within a few days.

Generator (hours run): 10,5 hours

Solar— SOC % – Turned off Gen : 100%, Turned on Gen : 69%

Diesel Reading – 75%

Propane Reading – 85%

Ethanol Free Gasoline – 7 gallons

Water (Auxillary tank) – Not in use

Water (Static tank) – 470 gallons

Water differential (Static tank) – 10 gallons

Gallons transferred: Not Applicable

Water in GreenHab – 264 gallons

Static to Loft Pump used – yes

Water Meter: 0141620,0

Toilet tank emptied: Yes

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Not used

Hours: 53,0 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Not used

Hours: 77,3 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Spirit rover used: Not Used

Hours: 78,4 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Nothing to report.

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

Hab Car used and why, where? Not used.

Staff Crew Car used and why, where? Off Campus

General notes and comments:

Summary of internet: Still on the fixing with BDM CapCom.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Noticed a disconnected tarp on
the connector between the Green Hab and the Science Dome
during this morning’s EVA. We have replaced it rapidly,
but it still needs fixation, especially considering the
wind. We will tackle this issue on the next EVA.

Summary of Green Hab operations: Nothing to report.

Summary of Science Dome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing more to report.

Operations Report – March 1st

Crew 206 Operations Report 01 – MAR – 2019

SOL: 5

Name of person filing report: Gaspard Thieulin

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and
alternator need replacement on generator. A contractor has
been arranged to make these repairs within a few days.

Generator (hours run): 10,5 hours

Solar— SOC % – Turned off Gen : 100%, Turned on Gen : 69%

Diesel Reading – 75%

Propane Reading – 85%

Ethanol Free Gasoline – 7 gallons

Water (Auxillary tank) – Not in use

Water (Static tank) – 470 gallons

Water differential (Static tank) – 10 gallons

Gallons transferred: Not Applicable

Water in GreenHab – 264 gallons

Static to Loft Pump used – yes

Water Meter: 0141620,0 gallons

Toilet tank emptied: Yes

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Not used

Hours: 53,0 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Not used

Hours: 77,3 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Spirit rover used: Not Used

Hours: 78,4 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Nothing to report.

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

Hab Car used and why, where? Not used.

Staff Crew Car used and why, where? Off Campus

General notes and comments:

Summary of internet: Still on the fixing with BDM CapCom.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Noticed a disconnected tarp on
the connector between the Green Hab and the Science Dome
during this morning’s EVA. We have replaced it rapidly,
but it still needs fixation, especially considering the
wind. We will tackle this issue on the next EVA.

Summary of Green Hab operations: Nothing to report.

Summary of Science Dome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing more to report.

Astronomy Report – March 1st

Astronomy Report, Sol 5

Name: Aurélien Mure

Crew: 206
Date: 03/01/2019

MUSK OBSERVARORY

Features observed : None

Problem encountered : Can you tell me if it is normal that no feature
can be focused these days ?

ROBOTIC OBSERVATORY

Object observed : M106

I did not attach the image I did from the observation because I did not
succeed to stack the images the right way to give a good final image.
I’ll continue to work on it.

Problems encountered : None

EVA Report – March 1st

Crew 206 EVA Report 01-MAR-2019
EVA #4
5
Author: Benjamin Auzou

Purpose of EVA:
The EVA is built to continue our exploration of different regions of
MDRS area. This one will be done without vehicule and so with only three
astronauts. We will climb to Hab Ridge by the 1103 path. Once up there,
the purpose will be to go through Hab Ridge Road and Sagan Street
intersection, to walk along the ridge and to arrive at Hab Ridge. On our
way we will make a stop to take picture of our home from Hab view North.
During the EVA we will also walk around the station in order to spot
degradations after the snowstorm.

Start time: 9:00 am

End time: 10:52 am

Narrative: Today, we went pretty fast on our EVA preparation like
yesterday, so we left the hab at 0900, instead of 0915, as it was
planned. Starting from tomorrow we will reconsider our preparation time
to leave the Hab at planned time.
That was our first EVA with an odd number of members, that was
unsettling for the preparation time, since we are usually working with
pairs (Team Blue and Red) for security checks. For the rest of the EVA
that change was less disruptive because of the absence of rover travels
and the good cohesion of our EVA team.
After the usual checks (Water tank for Team Red, composed of me, and
rovers states for Team Blue, composed of Cerise and Gaspard), we walked
around the MDRS campus in order to spot eventual degradations. The only
thing we spotted was a piece of tarp from the corridors that has been
moved by the wind. At 0907, all checks had been done and we started our
walk to Hab Ridge.
Looking for the road 1103, we stumbled upon a large black tarp (see
attached picture) at the coordinates 518418/4251029 (read on the GPS).
On the climb to Hab Ridge, the path was still muddy and slippery, even
if the surface seemed dry. However we managed to pay attention to each
other and walked safely toward the ridge.
After a couple minutes being amazed by the geologic diversity of the
landscape, we continued our walk to our goal : Hab View North.
We arrived at this point at 1002. We spent 10 minutes here enjoying the
view and taking some pictures of the Hab from the ridge.
Then we decided to return to the Hab following our steps in the opposite
direction instead of trying to go directly down the ridge. Safety first !
We arrived at the station and entered the airlock at 1047.

Destination: Hab Ridge Road / Sagan Street , Hab ridge, Hab view north

Coordinates: 517900/4251200 (Hab Ridge Road /Sagan Street),
517700/4250800 (Hab ridge), 517900/4250900 (Hab view North)
EVA Participants: Benjamin Auzou (COMMS, EVA leader), Gaspard Thieulin
(ENG), Cerise Cuny (HSO).
Roads and routs per MDRS Map: Road 1103 , Hab Ridge Road

Mode of travel: Walking

Journalist Report – March 1st

Crew 206 – 03/01/2019

Sol 5

Author: Auzou Benjamin, Journalist

"Music, Music"

The goal of the EVA of the morning was to climb to Hab Ridge to reach
Hab View North and enjoy a nice view on the station. With Cerise and
Gaspard, we faced the muddy slopes during an hour before reaching our goal.
The landscape up there was fantastic; every quarter turn I made around
me was the discovery of a new landscape, a new universe. I was
travelling from round and eroded rocks to sharp buttes, from an abrupt
plateau that reminded me of Monument Valley to a snowy mountain. That
was like if this place had been a draft for the creation of geologic
richness and diversity of this planet.

Since the beginning of our mission, music has a precious role, in
musical ambiance for life moments, sport sessions or experiments, to
relax before going to sleep or just to spend time and have fun. Everyone
shares its favorite playlist, its favorite music style.

We offer you a musical selection, composed of six songs, one for each of
us, that we listened, discovered or rediscovered during the beginning of
the mission.

-Come Together, The Beatles
"This music has the power to give me back energy even in fatigue
moments, so that it’s difficult for me to not singing. It’s a perfect
song to resist fatigue or boredom moments, but really for my crewmates’
ears."

-Papa loves Mambo, Perry Como
"Listened during reports writing, this song was relaxing and catchy for
the crew, that was a good moment for me and the rest of the team!"

-Mr Brightside, The Killers
"This song represents the beginning of my adventure in loving english
rock, as it is the first song I actually recall loving. I remember with
remarkable details the first time I heard it in my brother’s room when I
was 6. Mr Brightside was played randomly yesterday evening in the
playlist of an other crew member."

– Contact, Daft Punk
"When I listen to this music, it feels like I am with the astronauts, on
their way to space. It gives me goose bumps listening to the beats and
to the rocket taking off, but also thinking of what they’ve achieved."

– Sleepless Nights, FM Attack
"This title moves me to another place temporarily, on a sunny beach or
on the shore of a calm and peaceful lake. It helps me concentrate, find
energy and motivation to work on any task. Even for a short time, it
helps me continue for a while."

– Banana Pancakes, Jack Johnson
"For some reason, this song never fails to remind me of our little trips
to the beach in family when I was a kid. Playing it during the mission,
so far away from those moments, made me feel almost at home again, and
as always, it also filled me with happiness and good memories."

Greenhab Report – March 1st

Crew 206 Green Hab Report – 01-MAR-2019

GreenHab Officer: Norbert Pouzin

Environmental control: Heating and Cooling w/ ambient air (8 hrs)

Shade Cloth (40%)

Average temperatures:

Low: 17.5°C

High: 29.8°C

Hours of supplemental light: 0 Hours

Daily water usage for crops: 6.4 gallons

Water in Blue Tank: 262.7 gallons

Crops watered at 0900 hours (3.2 gallons) and 1830 hours (3.2 gallons):

TIME NOTED TEMPERATURE(° C ) / HUMIDITY(%)

1) 09.00 AM 20.1° C / 32 %
2) 12.00 AM 25.7° C / 25 %
3) 15:00 PM 28.6° C / 20 %
4) 18:30 PM 21.3° C / 30 %

Changes to crops: I removed the tomatoes and prepared the soil for the
next seeds.
I harvested carrots.
Radishes are germinating.
Red lettuces have some problem growing up, and are still not red.

Narrative:
The monitoring of the fan that I made yesterday worked well today, and
helped control the heat, even if it is still really hot when the sun
shines. Maybe, in the next weeks, I will put on the cooling system.
Indeed, it appeared that for a moment the temperature increased even
with a fan at max speed.

I also discovered that the "Crew 205 ‘s mystery seeds" were… radishes
! So I pulled out some radishes from this planter to allow the biggest
ones to grow.

Radishes from my experiment have a great germination : more than 75% of
the planted seeds have germinated, with an average of 2.6 cm of height.

Harvest: 9.0 g of basil, 2.1 g of parsley, 31.4 g of radish, 2.3 g of
lemon basil

Support/supplies needed: None.

Sol Summary – March 1st

Crew 206 March 1st 2019
Sol 5

Summary Title : Debating Day

Author : Jeremy Auclair

Mission Status : All nominal

Sol Activity Summary : Waking up was harder this morning, we all felt
quite tired. Muscle aches have gone but sport was hard nonetheless.
Benjamin, Cerise and Gaspard went by foot on top of the Hab ridge for
their EVA. They loved the view and did a wonderfull EVA, but they were
kind of hot around 10:00 am. Aurélien, Norbert and I worked on various
projects in the Hab. Norbert worked in the GreenHab, Aurélien prepared
some lunch while I tried to fix some issues with the TELEOP experiment.
There’s some improvement, but it’s not yet done. We ate a light lunch,
and after that we talked for a while. We are all feeling some effects of
isolation and confinement, but we are doing fine. Some of us took a nap
during the afternoon while others worked on their experiment. Aurélien
prepared some bread for diner, it looks great. As the sun sets we are
more and more tired, tonight will be a long night for most of us.

Look ahead plan : Tomorrow morning 4 of the crew will go on an EVA to
test the simulated space suit, they are impatient. In the afternoon I
will try to make TELEOP work, and tomorrow night will be movie night
before the rest day.

Anomalies in work : None

Weather : Sunny in the morning, partly cloudy in the afternoon, wind
increased after 04:00 pm.

Crew physical status : Fine

EVA : Hab Ridge

Reports to be filed : GreenHab, Journalist, Operation, EVA Report, EVA
Request, Astronomy

Support Requested : Lunar calendar (we want to take night time picture
of the sky)